OpenStudy (anonymous):

derive f(x) = (x+2)^(2/3) I know that you would use the chain rule but if someone could properly go over each step in deriving this it would be great since my teacher didn't really show us how to...

OpenStudy (anonymous):

thanks for the help in advance!

OpenStudy (anonymous):

\(\large f(x) = (x+2)^{2/3} \) \(\large f'(x) = \frac{2}{3}(x+2)^{2/3-1} \) \(\large f'(x) = \frac{2}{3}(x+2)^{-1/3} \) \(\large f'(x) = \frac{2}{3(x+2)^{1/3}}=\frac{2}{3\sqrt[3]{x+2}} \)

OpenStudy (anonymous):

thanks a bunch man! I really appreciate it!

OpenStudy (anonymous):

np... :)